OpenStudy (sydthekid913):

What is the y-intercept of the equation 3y + 2x = 12?

OpenStudy (anonymous):

Let x = 0, and solve for y.

OpenStudy (sydthekid913):

ok so it would be 0+2x=12?

OpenStudy (anonymous):

let x = 0

OpenStudy (sydthekid913):

O x

OpenStudy (sydthekid913):

so it would b 3y + 0?

OpenStudy (anonymous):

\[3y+2(0) = 12\]

OpenStudy (sydthekid913):

ok now what?

OpenStudy (anonymous):

Solve for y.

OpenStudy (sydthekid913):

Is it 4?

OpenStudy (anonymous):

yes

OpenStudy (sydthekid913):

so my final answer would be (0,4)?

OpenStudy (anonymous):

yes

OpenStudy (sydthekid913):

kool!

OpenStudy (sydthekid913):

Do I always substitute x for 0?

OpenStudy (anonymous):

Only when you're finding the y - intercept. If it was x - intercept, then you let y = 0.

OpenStudy (sydthekid913):

So it is always 0?

OpenStudy (anonymous):

Yes

OpenStudy (sydthekid913):

what would I do in a problem like this? A high school basketball team has a budget specifically for towels and extra basketballs. A towel costs $6 and a basketball costs $30. Assume the budget is $300 and that 25 towels are needed for the team for the entire season. Use a standard form equation to determine how many basketballs may be purchased by the team.

OpenStudy (triciaal):

6t + 30b = 300 t = # towels and b = # balls 25 towels needed so let t = 25 6*25 + 30b = 300 solve for b

OpenStudy (triciaal):

did you get 5 balls?